If I was to build stairs from my porch that is 6 feet high and and 7 feet from my house how many feet will my steps have to be
Answer by stanbon(75887)

If I was to build stairs from my porch that is 6 feet high and and 7 feet from my house how many feet will my steps have.
----
7x = 6
x = 6/7 ft = 10.288 inches
----
Each rise is 10.288 inches, if the depth of each step is 1 ft.
----------------
Each rise could be less that 10.288" if the depth of each
step is less than 1 ft.
Ex: If each depth is 8" = (2/3)'
then you need 6/(2/3) = 18/2 = 9 risers.
Then each riser would be 7'/9 = 84"/9 = 9 1/3 inches
======================
Cheers,
Stan H.